Circular No. 8040

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002kf S. Brady, Hudson, NH, reports the discovery of an apparent supernova

(mag ∼ 15.9) on an unfiltered CCD image taken with a 0.25-m automated telescope on Dec. 27.1 UT. The new object is located at α = 6^{h}37^{m}15^{s}.31, δ = +49^{o}51′10′′.2 (equinox 2000.0), which is ∼ 13′′.6 west and 10′′.6 north of the nucleus of the apparent host galaxy. SN 2002kf was confirmed on an unfiltered CCD frame taken by M. Schwartz with the Tenagra II 0.81-m telescope Dec. 28.4, but it does not appear on second Palomar Sky Survey red or blue plates.

COMET C/2002 V1 (NEAT) Ephemeris from orbital elements on MPEC 2002-Y60 (the magnitudes

below assume H_{15} = 6.5, though any prediction of the comet’s brightness close to perihelion is necessarily highly uncertain):

2002/03 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 22 1^{h}55.^{m}52 +14^{o}39.^{′}8 0.798 1.561 122.^{o}0 32.^{o}3 8.9
27 1 27.80 +13 46.5 0.797 1.466 110.3 39.0 8.5
Jan. 1 1 01.99 +12 50.3 0.807 1.367  99.1 45.3 8.1
6 0 38.50 +11 54.8 0.825 1.265  88.3 51.0 7.6
11 0 17.28 +11 02.1 0.847 1.159  78.2 56.1 7.1
16 23 58.01 +10 12.1 0.871 1.049  68.6 60.8 6.5
21 23 40.13 + 9 23.0 0.894 0.933  59.3 65.1 5.8
26 23 22.94 + 8 30.8 0.916 0.810  50.3 69.3 4.9
31 23 05.57 + 7 28.0 0.934 0.679  41.3 73.4 3.8
Feb. 5 22 46.83 + 6 00.5 0.947 0.536  32.1 77.9 2.3
10 22 24.90 + 3 35.5 0.957 0.377  22.3 83.3
15 21 56.89 − 1 38.4 0.970 0.194  11.3 89.6
20 21 53.19 −17 11.6 1.014 0.134  7.5 75.3

Visual m_{1} and coma-diameter estimates: Nov. 25.16 UT, 13.2, 1′.0 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 28.71, 12.1, 1′.5 (K. Yoshimoto, Yamaguchi, Japan, 0.25-m refl.); Dec. 4.60, 11.4, 2′.9 (Y. Nagai, Yamanashi, Japan, 0.32-m refl.); 7.89, 10.8, 4′ (J. M. San Juan, Toledo, Spain, 0.45-m refl.); 12.71, 10.3, 4′.0 (K. Yoshimoto, Yamaguchi, Japan, 0.25-m reflector); 23.69, 9.2, 8′ (K. Sárneczky, Piszkéstető, Hungary, 20×60 binoculars); 29.73, 8.0, 16′ (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars).

2002 December 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green